OpenStudy (waheguru):

what is the differnce between -16.1 and -14.7

OpenStudy (anonymous):

the answer is -1.4

OpenStudy (waheguru):

so the differnce is -1.4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-16.1 - (-14.7) = -16.1+ 14.7 = -1.4

OpenStudy (anonymous):

\[-112*\sqrt{3}*x^4*y^3*\sqrt{y}\]